OpenStudy (anonymous):

Put the equation in slope-intercept form and identify the slope of the line. 6x + 9y = 3 @TheStudyOwl A) -3/2 B)-2/3 C)-1/3 D)none of the above

OpenStudy (anonymous):

6x+9y=3 9y=-6x+3 y=-2/3x+1/3 the answer is -2/3 which is B..

OpenStudy (anonymous):

@vanthando Thanks so much, this test is a bit diffcult for me can you help me with two more questions?

OpenStudy (anonymous):

Given the domain values { -2, 0, 2 }, find the range values for the function y = 5x + 1. { -10, 1 10 } { -9, 1, 11 } { -11, 1, 11 } none of the above